madisonlacroix:

A water park charges a rental fee plus an hourly fee for inflatable inner tubes. The total costs of 3, 4, and 5 hours are $8.00, $8.75, and $9.50, respectively. Make a table or draw a graph to determine the rental fee and the hourly fee the water park charges. What is the cost of the hourly fee? What is the cost of the rental fee?

XioGonz:

I can't show you the exact work but I'll give an example of the steps. x = the rental fee x + 1.50*6 = 15 x + 9 = 15 x = 15 - 9 x = $6